The Lakeside Coffee Shop morning customer load follows a normal distribution. The mean number of customers is 60 and the standard deviation is 15. Determine the probability that the number of customers tomorrow will be 50. Include evidence of your work by typing out your solution fully.



1
Expert's answer
2021-08-04T14:41:01-0400

Let "X=" the number of customers: "X\sim N(\mu, \sigma^2)."


"f(x, \mu, \sigma)=\dfrac{1}{\sqrt{2\pi}\sigma}e^{-(x-\mu)^2/(2\sigma^2)}, -\infin<x<\infin"

"P(X=50)=P(50\leq X\leq50)=\displaystyle\int_{50}^{50}f(x)dx=0"

The probability that the number of customers tomorrow will be 50 is equal to 0.